OpenStudy (anonymous):

5/3x=-20 solve show work

OpenStudy (anonymous):

3x=25 x=25/3

OpenStudy (anonymous):

im not sure of ur answer

OpenStudy (anonymous):

5/3x = -20 3 * (5/3x) = 3 * -20 5x = -60 x = -12 is mine

OpenStudy (anonymous):

In easy to follow steps: 3x - 5 = 20 3x = 20 +5 (you add 5 to both sides. Adding to the left side, removes the -5) 3x = 25 x = 25/3 (divide both sides by 3) x = 8.3 The goal in this kind of equation is to solve for x. What you do to one side, you can always do to the other and you must do it to both at the same time.

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20090901170058AANpWGg <--------- go there

OpenStudy (anonymous):

correct but its a fraction not 3x-5

OpenStudy (anonymous):

Isolate x by multiplying by 3/5 x=-60/5 x=-12

OpenStudy (anonymous):

My answer is -(1/12) or -12. If it is 5 / (3X) = -20, then 5 = -20 * (3X) 5 = -60 * X -(1/12) = X If it is (5/3) * X = -20, then X = -20 * (3 / 5) X = - 12 So the answers could be -(1/12) or -12.

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

medal? please

OpenStudy (anonymous):

can i please get a medal